log2(x - 2) + 5 = 8 - log2 4
Answer by mukhopadhyay(490)

log2(x - 2) + 5 = 8 - log2 4
=> log(base2) (x-2) + log(base2) 4 = 8-5
=>log(base2) 4(x-2) = 3
=> 4(x-2) = 2^3
=> (x-2) = 8/4 = 2
=> x=4
Answer: x=4